2 wheels that both have numbers 1-8 are spun. Assuming the wheels are independent and the outcomes are equally likely, determine the probability of obtaining even numbers on both wheels.
Answer by stanbon(75887)

2 wheels that both have numbers 1-8 are spun. Assuming the wheels are independent and the outcomes are equally likely, determine the probability of obtaining even numbers on both wheels.
----
P(even on both wheels) = (4/8)(4/8) = 1/4
-----------------
Cheers,
Stan H.
===============